You turned the knob of a 196.3 cm by 109.1 cm door (with a mass of 9.88 kg) and pushed it with a perpendicular force of 9.46N to open it. What is the angular acceleration of the door (in rad/s2)?


1
Expert's answer
2021-11-29T11:44:36-0500

"\tau=I\epsilon \to F\cdot d=I\epsilon \to \epsilon=F\cdot d/I="


"=F\cdot d/(ma^2/3)=3F\cdot d/(ma^2)="


"=3\cdot9.46\cdot1.091/(9.88\cdot 1.091^2)=2.63\ (rad/s^2)" . Answer